A chemist has a 100 mL beaker of a 30% solution.
a. Write an equation for the concentration of the solution after adding x mL of pure water.
b. Use that equation to determine how many mL of water should be added to obtain a 7% solution. Round your answer to 1 decimal place.

Answer by stanbon(75887)

A chemist has a 100 mL beaker of a 30% solution.
a. Write an equation for the concentration of the solution after adding x mL of pure water.
--------
Equation:
active + active = active
0.30*100 + 0*x = c(100+x)
============================
b. Use that equation to determine how many mL of water should be added to obtain a 7% solution. Round your answer to 1 decimal place.
------
30*100 + 0*x = 0.07(100 + x)
30*100 + 0 = 7*100 + 7x
7x = 21*100
x = 300 mL (amt of water to add)
=============
Cheers,
Stan H.